# Insight-HXMT View of the BHC Swift J1727.8-1613 during its outburst in 2023

### v2

@@ 2.3 Data Analysis
We have downloaded the daily average light curve of MAXI/GSC [^5] http://maxi.riken.jp/star_data/J1727-162/J1727-162.html from the archive to study the outburst profile in addition to Insight-HXMT. Using the unbinned light curves from LE, ME, and HE modules, we have first produced $0.01$ sec time-binned light curves. Then the power density spectrum (PDS) was generated using the fast Fourier transformation (FFT) technique in the powspec task of the XRONOS package in the HEASoft software. Each observation’s data was subdivided into several intervals, each of which contains 8192 newbins. First, the PDS for each interval is generated and then they are averaged to make a resultant PDS. The PDS is normalized in a way that their integral can give the RMS-squared fractional variability. The white noise level is subtracted and a geometrical rebinning of -1.02 is used. For some data, we needed to use a geometrical rebinning of -1.05 to have a prominent QPO-like nature in the PDS. We fit the QPO using single or multiple Lorentzian model components in powspec to extract QPO properties like frequency ($\nu_{qpo}$), full-width at half maximum or FWHM, and normalization. For several observations, we have found the presence of single or multiple harmonics. We have also extracted their properties using the same fitting method. We have fitted the light curves of all the exposures of the listed 10 observations. We report them in the result section.
For the spectral analysis, we have used the same LE, ME, and HE modules and fitted the broadband data in the $2-150$ keV energy band. The best is achieved by using the combination of disk blackbody, power-law, gaussian, and pexrav models. For some observations, we needed to add a gabs model to achieve the best fit. For interstellar absorption, we have used the tbabs model. Since we simultaneously fit all three modules, we have included a constant to normalize the three resultant fittings. The following two model combinations are used: i) constant*tbabs*gabs*(diskbb + power-law + gaussian + pexrav), ii) constant*tbabs*(diskbb + power-law + gaussian + pexrav) (for which gabs was not required. The spectral analysis was done for the exposures are marked with ‘*’ in Table 2. We did not do the spectral analysis for all epochs as spectral variability was not present within such a small timescale (intraday) like the variability in the light curves. We report them in the result section.
@@ 3 Results
Using the timing and spectral analysis, we have studied the accretion flow properties of the very recent 2023 outburst of the BHC Swift J1727.8-1613. We studied the QPO properties of the source during the outburst as well as the spectral nature and radiation properties using Insight-HXMT data. We broadly discuss the results in the next two sections.
@@ 3.1 Temporal Properties
Here, we discuss the evolution of the flux and hardness ratio, and QPO properties during the first 10 observation IDs of the outburst.
@@ 3.1.1 Outburst Profile and Hardness Ratio
In panel (c), we show the variation of the hardness ratio, which is the ratio of the hard ($F_{H}$) $6-20$ keV to the soft ($F_{S}$) $2-6$ keV flux. At the start of the outburst, the HR was $\sim 0.5$ and then it gradually decreased to MJD 60235 as the outburst progressed. After that, the HR became almost constant, although decreasing very slowly to $\sim 0.025$ at the end of the considered data of the outburst on MJD 60350. The way HR is defined, when this value is on the higher side, the spectral nature of the source is hard. This is the general case when an outburst is starting. When this value is low, there is dominance of soft flux over hard flux and the spectral nature would be soft. Without studying the detailed spectral analysis, HR provides a quick idea about the spectral nature of an outburst. However, for this source, we notice that HR was a little higher at the onset of the outburst, where the hard flux was also comparable to the soft flux. However, after a few days, soft flux completely took over the hard flux, and the HR was also not high. Depending on the HR value, we propose that the source was already in the SIMS after the onset of the outburst. After MJD 60185, it made a transition into the SS. Since we are not analyzing the full duration of the outburst, it is not the scope of the paper to study the full spectral classification. We only define the spectral state of the source based on the duration of which data is used in this work.
@@ 3.1.2 Quasi Periodic Oscillations (QPOs)
As described in §2, we used the $0.01$ sec time-binned light curves from all three bands (LE, ME, and HE) and produced the PDS to study QPOs. We find that QPO nature was present in all the light curves in all three energy bands. In Figure 2, we show the variation of the QPO frequency ($\nu_{qpo}$) for the 10 different obs. IDs as stated before. We first checked the first obs. ID P0614338001. It has several exposure IDs (given in Table 2). While we fit every exposure of this obs ID, we found that the QPO frequency was changing even within this small period. This is the reason we fitted all the light curves of the exposures of the obs IDs that we have taken for analysis. We find that for every single obs ID, the $\nu_{qpo}$ showed significant variation. Figure 2(a), (b), (c) show the variations of the $\nu_{qpo}$ with time for 10 different obs IDs for LE, ME, and HE respectively.
Evolution of QPO frequency in different observation ID.s
@@
In Figure 3, we show the evolution of the $\nu_{qpo}$ for the total duration of our analysis period. In (a), (b), and (c) panels the variations are for LE, ME, and HE respectively. We find that the QPO frequencies were $\sim 0.23\pm 0.008$, $0.24\pm 0.002$, and $0.24\pm 0.002$ Hz for LE, ME, and HE on the starting day of our analysis. After this day, the $\nu_{qpo}$ shows a rapid increase with time till $\sim$ MJD 60187 when $\nu_{qpo}$ was $\sim 1.2$Hz for all the three bands. After this date, it increased very slowly till MJD $\sim 60190$ when $\nu_{qpo}$ was $\sim 1.45$. Then it decreased for a very short period until it rapidly increased to $1.5$Hz again in all three bands. After that, it showed varying nature till MJD $\sim 60197$, and then it finally increased again. We have discussed the possible reason behind this nature in the next section.
Using the PDS fitting, we extracted some information (e.g., full-width at half maximum (FWHM), Normalization (LN)) of QPOs. We also extracted source and background count rates for the exposures. Using these estimates, we calculated the $Q$-value and RMS that represent the sharpness of the QPO and the fractional variability in the PDS, respectively. These values are listed in Table 3 in columns 5–7 ($Q$-value) and 8–10 (RMS) for LE, ME, and HE respectively.
Evolution of RMS (%) with QPO frequency and time
Refer to captionRefer to caption
Figure 4: 3-D representation of the variation of $\nu_{qpo}$ and (%)RMS with time (MJD) for (a) LE, (b) ME, and (c) HE.
In Figure 4, we show a 3-D representation of the evolution of QPO and RMS with time. In the LE band (Fig. 4a), the RMS was higher at the start of the analysis period and it then gradually decreased. In the ME band (Fig. 4b), it increased from the onset of the analysis period for some time and then again decreased. The HE band (Fig. 4c) shows a similar trend as in LE band. Since the QPO nature has already been reported in the recent paper by Yu et al. (2024), we are not focusing on the classification of the QPO nature. However, by the values of both $Q$-factor and RMS, it can be said that the QPO is type-C in nature for all the exposures reported here.
@@ 3.2 Spectral Properties
We first fitted the data with phenomenological diskbb and power-law models. We used the multiplicative tbabs model to account for the interstellar absorption. Our model read as constant*tbabs(diskbb + powerlaw) in XSPEC. The constant is taken to normalize between three different instruments in LE, ME, and HE. We found that the fit was not statistically acceptable. There was a clear sign of the Gaussian feature in the unfolded spectrum around $5-6$ keV. Also, there was a hump-like feature $\sim 20$ keV. Besides these features, the absorption feature was noticed below 2 keV. Thus, we refitted the data using the model combination as constant*tbabs*gabs(diskbb + power-law + ga + pexrav). Here, the gabs model takes care of the low energy absorption feature, while the ga and pexrav models take care of the contribution from the Gaussian and hump-like nature, as mentioned above. In Fig. 5a, we show the best-fitted unfolded spectrum using this model combination for the observation ID P061433800501 (Exposure ID: P061433800501-20230902-01-01), where $\chi^{2}/DOF\sim 1$. After some exposure, we found that the gabs model was not needed anymore to achieve the best fit. Then, the model combination in XSPEC reads as constant*tbabs(diskbb + power-law + ga + pexrav). In Figure 5b, we show an unfolded best-fitted spectrum using the later model combination. This is for the observation ID. P061433800601 (exposure ID: P061433800601-20230904-01-01), for which $\chi^{2}/DOF\sim 0.9$.
Model Fitted Parameters
@@
Figure 6: Variation of spectrally fitted properties (a) hydrogen column density ($n_{H}$), (b) inner-disk temperature ($T_{in}$), (c) diskbb normalization, (d) Photon Index ($\Gamma$), (e) reflection fraction (Rref), and (f) cutoff energy (Ecut) with time.
@@ 3.3 Correlation between spectral and timing properties
The QPO frequency can be produced from the oscillation of the Compton corona, which can also be the origin of the hard power-law photons that extend the spectrum to higher energies. Therefore, these two features might be interlinked or correlated. It has been previously seen that the QPO frequency ($\nu_{qpo}$) and $\Gamma$ show a positive correlation for some of the sources during their active outbursting phase (Vignarca et al. 2003; Shaposhnikov & Titarchuk 2009; Stiele et al. 2013). We also tried to find out if there are any correlations between these two properties. In Figure 7, we show the variations of the photon index with the QPO frequency for the three energy bands. Considering these points, we have found a positive correlation between $\Gamma$ and $\nu_{qpo}$. Using both the Pearson-Linear and Spearman-Rank correlation methods, we found correlation coefficients for the three energy bands as shown in the table below.
@@ 4.1 Propagation of Shock
Quasi-periodic oscillation is one of the most important and common phenomena for stellar mass black holes. For this newly discovered source, we analyzed a total of 342 light curves (114 for LE, ME, and HE each). However, in some of the exposures, light curves were not produced properly. There were a total of 92 exposures which commonly produced LE, ME, and HE light curves. We list those data in Table 3. Finally, we analyzed 276 lightcurves. We found the presence of QPO in all of these light curves, and they showed rapid evolution in frequency in this short analysis period. The QPO frequency ($\nu_{qpo}$) showed an increase in frequency within even a single day. The evolution of the QPO frequency has been discussed broadly in the result section. It showed an increasing trend from the start of our analysis. From the variation of QPO frequency, (%)RMS, and $Q$-factor, the nature of the QPO is designated to be type-C. Although the QPOs are extensively studied in the literature, their origin requires in depth modeling.
Here, we would like to focus on the physical scenario that explains the origin of QPOs due to the shock instabilities in advective flows around BHs. Accretion onto BH is transonic flow with possibility of multiple sonic points (Chakrabarti 1989). The matter from the companion need not necessarily be only Keplerian. There could be a supply of matter which has angular momentum distribution deviated from the Keplerian one. This is the sub-Keplerian component. Such a lower angular momentum component of infalling matter gets accreted in a free-fall timescale. Due to the tug of war between the gravitational force and centrifugal force, this matter could virtually stop at some distance from the BH and undergo shock transition, creating a post-shock region. Depending on the flow properties, the standing shocks forms (Chakrabarti 1989; Singh et al. 2022 and references therein). This approach can well describe the observed spectral and temporal properties of BHs (Debnath et al. 2014; Mondal et al. 2014, Chatterjee et al. 2020, 2021, 2023). According to Molteni et al. (1996), the shock can be oscillatory due to the presence of cooling. This was later shown in simulation by Garain et al. (2014). During the oscillation when the cooling timescale due to the Comptonization process and the heating timescale matches, QPOs originate (see Chakrabarti et al. 2015). Additionally, it has also been observed by the authors that once the QPO sets in, it gets locked for a week or more depending on the above condition. From the expressions of infall and cooling timescales, we get that the QPO frequency ($\nu_{qpo}$) can be estimated using the relation (Molteni et al. 1996; Chakrabarti & Manickam 2000; Chakrabarti et al. 2004),
Where $c$, $G$, $M_{BH}$, $X_{s}$, and $R$ are the speed of light, gravitational constant, the mass of the BH, location of the shock, and the ratio of matter densities in post-shock to pre-shock regions ($\rho_{+}/\rho_{-}$) respectively. From our timing analysis, we already extracted the information about QPO frequency ($\nu_{qpo}$). Using the above relation, we estimated the shock location during the outburst. According to Chakrabarti et al. (2005), $X_{s}$ can be located anywhere above 10 $r_{s}$ depending on the flow parameters producing shocks. The shock forms far away $\sim 1000~{}r_{s}$ when the spectral nature of the ongoing outburst is hard and decreases gradually in the progressive days as cooling increases (Mondal et al. 2015). Using Eq. 1, we found that the shock was far away from the BH at the start of the outburst (see Figure 8b). It then significantly decreased as the QPO frequency increased rapidly during the initial few days. Close to MJD 60185, when the $\nu_{qpo}$ stopped increasing, and the shock location started to decrease very slowly. The values of the shock location along with QPO frequency are given in Table 3.
@@
Using this relation, we have estimated the velocity of the movement of the shock. The estimated shock velocities are given in Figure 8c, which is a few tens of m/s in accord with the previous estimations (Chakrabarti et al. 2005; Mondal et al. 2015). We find that the shock velocity has varied within this period. At the beginning, it attained higher values when the $X_{s}$ was also larger, however decreasing significantly afterward. It slowed down as the shock moved inwards, along with some fluctuations. This can be due to the limitation of assuming constant velocity of the shock propagation.
4.2 Evolution of the physical configuration and spectral nature

### v3

@@ 2.1 Data Selection
Here, we note that, in general, the exposure is the time span of observation within an observation ID. The Insight-HXMT has several observation IDs for this source, each of which lasts for a different amount of time, which is the exposure time. In Table 1, one can see that all the observation IDs have different exposures, with some of them having a time gap between them. When reducing the data, we found that inside every observation ID, there were several sub-IDs, the name of which starts with that particular ID. The total exposure of an entire observation ID was cut into several sub-IDs with different exposure times but with no time gap between them. This can be seen in Table 3 (columns 3 and 5). GTIs are also different. They can be separated in time between the consecutive ones. Here, this is not the case. Thus, to differentiate between them, we called these as ‘exposure IDs’, and used this phrase throughout the paper. Within 1 observation ID (like the first one in Table 1), there are several exposure IDs (like 12 exposure IDs in the 1st column of Table 3). This has not been defined like this before. To keep things easy, we have taken this approach.
@@ 2.3 Data Analysis
We have downloaded the daily average light curve of MAXI/GSC [^5] http://maxi.riken.jp/star_data/J1727-162/J1727-162.html from the archive to study the outburst profile in addition to Insight-HXMT. Using the unbinned light curves from LE, ME, and HE modules, we have first produced $0.01$ sec time-binned light curves. Then the power density spectrum (PDS) was generated using the fast Fourier transformation (FFT) technique in the powspec task of the XRONOS package in the HEASoft software. Each observation’s data was subdivided into several intervals, each of which contains 8192 newbins. First, the PDS for each interval is generated and then they are averaged to make a resultant PDS. The PDS is normalized in a way that their integral can give the RMS-squared fractional variability. The white noise level is subtracted by using a normalization value of $-2$ in powspec and a geometrical rebinning of -1.02 is used. For some data, we needed to use a geometrical rebinning of -1.05 to have a prominent QPO-like nature in the PDS. We model the PDS with Lorentzian model for both the fundamental QPO and the harmonic components in powspec. The parameters from the fitted Lorentzian model for the fundamental QPO help us obtain the frequency ($\nu_{qpo}$), the full-width at half maximum or FWHM, and the normalization. For several observations, we have found the presence of single or multiple harmonics. We have also extracted their properties using the same fitting method. We have fitted the light curves of all the exposures of the listed 10 observations. We report them in the result section.
For the spectral analysis, we have used the same LE, ME, and HE modules and fitted the broadband data in the $2-150$ keV energy band. The best is achieved by using the combination of disk blackbody, power-law, gaussian, and pexrav models. For some observations, we needed to add a gabs model to achieve the best fit. For interstellar absorption, we have used the tbabs model. Since we simultaneously fit all three modules, we have included a constant to normalize the three resultant fittings. The following two model combinations are used: i) constant*tbabs*gabs*(diskbb + power-law + gaussian + pexrav), ii) constant*tbabs*(diskbb + power-law + gaussian + pexrav) (for which gabs was not required. The spectral analysis was done for the exposures are marked with ‘*’ in Table 3. We did not perform the spectral analysis for all epochs as we did not see change in spectral features over such a small time gap unlike the timing properties, which showed variations within a day. We report them in the result section.
@@ 3 Results
Using the timing and spectral analysis presented above, we have studied the accretion flow properties of the very recent 2023 outburst of the BHC Swift J1727.8-1613. We studied the QPO properties of the source during the outburst as well as the spectral nature and radiation properties using Insight-HXMT data. We broadly discuss the results in the next two sections.
@@ 3.1.1 Outburst Profile and Hardness Ratio
In panel (c), we show the variation of the hardness ratio, which is the ratio of the hard ($F_{H}$) $6-20$ keV to the soft ($F_{S}$) $2-6$ keV flux. At the start of the outburst, the HR was $\sim 0.5$ and then it gradually decreased till MJD 60235 as the outburst progressed. After that, the HR became almost constant, although decreasing very slowly to $\sim 0.025$ at the end of the considered data of the outburst on MJD 60350. The way HR is defined, when this value is on the higher side, the spectral nature of the source is hard. This is the general case when an outburst is starting. When this value is low, there is dominance of soft flux over hard flux and the spectral nature would be soft. Without studying the detailed spectral analysis, HR provides a quick idea about the spectral nature of an outburst. Depending on the HR value, we propose that the source was already in the intermediate state after the onset of the outburst. After MJD 60185, it made a transition into the SS.
@@ 3.1.2 Quasi Periodic Oscillations (QPOs)
Figure 2: Model fitted PDS using $0.01$ sec time-binned HE light curve from the observation ID P0614338004 (Exposure ID P061433800401-20230831-01-01).
@@
As described in §2, we used the $0.01$ sec time-binned light curves from all three bands (LE, ME, and HE) and produced the PDS to study QPOs. We find that QPO nature was present in all the light curves in all three energy bands. In Figure 2, we show a model fitted PDS continuum in the $0.1-50$ Hz frequency range for $0.01$ sec time-binned HE light curve from the observation ID P0614338004 (Exposure ID P061433800401-20230831-01-01). In Figure 3, we show the variation of the QPO frequency ($\nu_{qpo}$) for the 10 different obs. IDs as stated before. We first checked the first obs. ID P0614338001. It has several exposure IDs (given in Table 3). While we fit every exposure of this obs ID, we found that the QPO frequency was changing even within this small period. This is the reason we fitted all the light curves of the exposures of the obs IDs that we have taken for analysis. We find that for every single obs ID, the $\nu_{qpo}$ showed significant variation. Figure 3(a-c) shows the variations of the $\nu_{qpo}$ with time for 10 different obs IDs for LE, ME, and HE respectively. It is interesting to observe that $\nu_{qpo}$ has changed in frequency so much within this short period, which suggests that there are some changes happening in the system in a very short timescale.
@@
In Figure 4, we show the evolution of $\nu_{qpo}$ for the total duration of our analysis period. In (a), (b), and (c) panels the variations are for LE, ME, and HE respectively. We find that the QPO frequencies were $0.23\pm 0.01$, $0.24\pm 0.01$, and $0.24\pm 0.01$ Hz for LE, ME, and HE on the starting day of our analysis. After this day, $\nu_{qpo}$ shows a rapid increase with time till $\sim$ MJD 60187 when $\nu_{qpo}$ was $1.16\pm 0.01$, $1.16\pm 0.01$, and $1.16\pm 0.01$ for LE, ME, and HE, respectively. After this date, it increased very slowly till MJD $\sim 60190$ when $\nu_{qpo}$ was $1.44\pm 0.01$, $1.44\pm 0.01$, and $1.50\pm 0.01$ for the three bands. Then it decreased for a very short period until it rapidly increased to $\sim 1.5$ Hz again in all three bands. After that, it increased and decreased in a narrow range of $\sim 1.0-1.5$ Hz till MJD $\sim 60197$, and then it finally increased again. We have discussed the possible reasons behind this nature in the next section.
Using the PDS fitting, we extracted the centroid frequency, the full-width at half maximum (FWHM) and the Normalization (LN)) of the QPOs. The ‘fplot’ task of the ‘ftools’ package in the HEASoft software produces plots of the light curves. Using this task, we plotted and saved each value of time and the corresponding count rate. Then we average the counts to find out the average count rate for that light curve. The same task is applied to extract both the source and background count rates for all the exposures. Then, using these estimates, we calculated the $Q$-value and RMS that represent the sharpness of the QPO and the fractional variability in the PDS, respectively. These values are listed in Table 4 in columns 5–7 ($Q$-value) and 8–10 (RMS) for LE, ME, and HE respectively.
Figure 5: Evolution of the QPO RMS (%) with time (MJD) for (a) LE, (b) ME, and (c) HE.
In Figure 5, we show the evolution of the QPO RMS (%) with time (MJD). In the LE band (Fig. 5a), the RMS was higher at the start of the analysis period and it then gradually decreased. In the ME band (Fig. 5b), it increased from the onset of the analysis period for some time and then decreased. The HE band (Fig. 5c) shows a similar trend as in LE band. Since the QPO nature has already been reported in the recent paper by Yu et al. (2024), we are not focusing on the classification of the QPO. However, by the values of both $Q$-factor and RMS, it can be said that the QPO is type-C in nature for all the exposures reported here.
@@ 3.2 Spectral Properties
We first fitted the data with the phenomenological diskbb and power-law models. We used the multiplicative tbabs model to account for the interstellar absorption. Our model read as constant*tbabs(diskbb + powerlaw) in XSPEC. The constant is taken to normalize between three different instruments in LE, ME, and HE. We found that the fit was not statistically acceptable. There was a clear sign of the Gaussian feature in the unfolded spectrum around $5-6$ keV. Also, there was a hump-like feature $\sim 20$ keV. Besides these features, the absorption feature was noticed below 2 keV. Thus, we refitted the data using the model combination as constant*tbabs*gabs(diskbb + power-law + ga + pexrav). Here, the gabs model takes care of the low energy absorption feature, while the gaussian and pexrav models take care of the contribution from the Gaussian and hump-like nature, as mentioned above. In Fig. 5a, we show the best-fitted unfolded spectrum using this model combination for the observation ID P061433800501 (Exposure ID: P061433800501-20230902-01-01), where $\chi^{2}/DOF\sim 1$. After some exposures, we found that the gabs model was not needed anymore to achieve the best fit. Then, the model combination in XSPEC reads as constant*tbabs(diskbb + power-law + ga + pexrav). In Figure 6b, we show an unfolded best-fitted spectrum using the later model combination. This is for the observation ID. P061433800601 (exposure ID: P061433800601-20230904-01-01), for which $\chi^{2}/DOF\sim 0.9$.
@@
Figure 7: Variation of spectrally fitted properties (a) hydrogen column density ($n_{H}$), (b) inner-disk temperature ($T_{in}$ in keV), (c) diskbb normalization, (d) Photon Index ($\Gamma$), (e) reflection fraction (Rref), and (f) cutoff energy (Ecut) with time.
@@ 3.3 Correlation between spectral and timing properties
It has been previously seen that the QPO frequency ($\nu_{qpo}$) and $\Gamma$ show a positive correlation for some of the sources during their active outbursting phase (Vignarca et al. 2003; Shaposhnikov & Titarchuk 2009; Stiele et al. 2013). We also tried to find out if there are any correlations between these two properties. In Figure 8, we show the variations of the photon index with the QPO frequency for the three energy bands. Considering these points, we have found a positive correlation between $\Gamma$ and $\nu_{qpo}$. Using both the Pearson-Linear and Spearman-Rank correlation methods, we found correlation coefficients for the three energy bands as shown in Table 2. In all of the three energy bands, we see moderately strong correlation when it comes to $\Gamma$ and $\nu_{qpo}$ as correlation coefficient is $\sim 0.5-0.8$. This shows the timing and spectral properties, the origin of which are the same geometric variation of the Compton corona, are moderately correlated in this source.
Table 2: Correlation coefficients between the QPO frequency ($\nu_{qpo}$) and photon index of power-law ($\Gamma$).
@@ 3.4 Propagation of Shock
In the TCAF paradigm, the oscillation of the shock is responsible for the origin of QPOs. When this shock front oscillates, it produces QPOs. The shock oscillation model has two types of timescales: the infall timescale ($t_{i}$) and the cooling timescale ($t_{c}$) and matching of these timescales gives the resonance condition responsible for the QPOs. The relation between the QPO frequency and shock location is given as (Molteni et al. 1996; Chakrabarti & Manickam 2000; Chakrabarti et al. 2004),
where, $c$, $G$, $M_{BH}$, $X_{s}$, and $R$ are the speed of light, the gravitational constant, the mass of the BH, location of the shock (in units of the Schwarzschild radius $r_{s}$), and the ratio of matter densities in post-shock to pre-shock regions ($\rho_{+}/\rho_{-}$) respectively. From our timing analysis, we extracted the information about the QPO frequency ($\nu_{qpo}$). Using the above relation, we estimated the shock location during the outburst. Using Eq. 1, we found that the shock was far away ($>500~{}r_{s},r_{s}=2GM_{BH}/c^{2}$) from the BH at the start of the outburst (see Figure 9b). It then shifted towards the black hole sharply as the QPO frequency increased rapidly during the initial few days ($\sim 250~{}r_{s}$). Close to MJD 60185, the shock location started to decrease very slowly and it reached to $\sim 100~{}r_{s}$ at the end of our analysis period. The values of the shock location are given in Table 5.
@@
Using this relation, we have estimated the velocity of the movement of the shock. The estimated shock velocities are given in Figure 9c, which is a few tens of m/s in agreement with the previous estimates (Chakrabarti et al. 2005; Mondal et al. 2015). We find that the shock velocity varies within this period. In the beginning, the shock moved with a speed of $\sim 50-100$ m s-1, when the $X_{s}$ was also larger. Then, after MJD $\sim$ 60183.7, it decreased a little bit with values $<50$ m s-1. Then it slowed down and was moving with a speed of $\sim 10-30$ m s-1 with some fluctuations and continued till the end of the observation period.
4 Discussion
Quasi-periodic oscillation is one of the most important and common phenomena for stellar mass black holes. For this newly discovered source, we analyzed a total of 342 light curves (114 for LE, ME, and HE each). However, in some of the exposures, light curves were not produced properly. There were a total of 92 exposures which commonly produced LE, ME, and HE light curves. We list those data in Table 3. Finally, we analyzed 276 lightcurves. We found the presence of QPO in all of these light curves, and they showed rapid evolution in frequency in this short analysis period. The QPO frequency ($\nu_{qpo}$) showed an increase in frequency within even a single day. The evolution of the QPO frequency has been discussed broadly in the result section. It showed an increasing trend from the start of our analysis. We found that the RMS (%) was very high ($>13\%$) for all the exposures. The $Q$-value was also high, though for some observations it was close to those QPOs that are types-A/B. However, the frequency and rms allow us to classify the QPOs as type-C, which are in agreement with Casella et al. (2005). Considering the values of RMS, $Q$-value and the type of QPO, we suggest that the source was in the HIMS. Recently, Yu et al. (2024) also reported the timing properties of this source for a longer period. The QPO frequency has varied in a much broader range in their report from 0.1 to 8 Hz. The authors also concluded the observed QPOs in the full period as type-C. During the 2009-10 outburst of the BHC XTE J1752-223 (Chatterjee et al. 2020), we found the existence of type-C QPOs that had $Q$-values comparable to this source. During the 2016-17 outburst of the BHC GRS 1716-249 (Chatterjee et al. 2021), type-C QPOs were found to be present with similar RMS (%) values. Other than these sources, Belloni et al. (2002), Casella et al. (2005), and the references therein can be seen for an analogy. Such similarities in observed QPO properties in different sources require physical explanations of common origin.
Here, we would like to focus on the physical scenario that explains the origin of QPOs due to the shock oscillations in advective flows around BHs. Accretion onto BH can be explained as transonic flow with possibility of multiple sonic points (Chakrabarti 1989). The matter from the companion need not necessarily be only Keplerian. There could be a supply of matter which has angular momentum distribution deviated from the Keplerian one. This is the sub-Keplerian component. Such a lower angular momentum component of infalling matter gets accreted in a free-fall timescale. Due to the tug of war between the gravitational force and centrifugal force, this matter could virtually stop at some distance from the BH and undergo shock transition, creating a post-shock region. Depending on the flow properties, the standing shocks forms (Chakrabarti 1989; Singh et al. 2022 and references therein). This approach can well describe the observed spectral and temporal properties of BHs (Debnath et al. 2014; Mondal et al. 2014, Chatterjee et al. 2020, 2021, 2023). According to Molteni et al. (1996), the shock can be oscillatory due to the presence of cooling. This was later shown in simulation by Garain et al. (2014). During the oscillation when the cooling timescale due to the Comptonization process and the heating timescale matches, QPOs originate (see Chakrabarti et al. 2015). Additionally, it has also been observed by the authors that once the QPO sets in, it gets locked for a week or more depending on the above condition.
According to Chakrabarti et al. (2005), $X_{s}$ can be located anywhere above 10 $r_{s}$ depending on the flow parameters producing shocks. The shock forms far away $\sim 1000~{}r_{s}$ when the spectral nature of the ongoing outburst is hard and decreases gradually in the progressive days as cooling increases (Mondal et al. 2015). When the cooling timescale ($t_{c}$) and infall timescale ($t_{i}$) becomes comparable to each other (Molteni et al. 1996), an oscillation could set in the shock boundary due to resonance. This could produce variations as we observe in the light curves. This might be the case for this source regarding the origin of the occurrence of the QPOs. Here, from this analysis, we have found that the shock was far away ($\sim 500~{}r_{s}$) when QPO frequency ($\nu_{qpo}$) was low ($\sim 0.2$ Hz). The disk was far away. As the disk moves inward, cooling increases with progressive days and QPO frequency increases, as shown in Eq. 1; following inverse scaling with $X_{s}$. This is the case that we observed here. After the start, as the outburst progressed, the cooling began to take place and $X_{s}$ got smaller in size. As a result, $\nu_{qpo}$ increased. After a few days, the shock location decreased very slowly, without changing its value much, the $\nu_{qpo}$ was also increasing very slowly. We also tried to observe how the shock moved, considering the constant velocity approach, as given in Eq. 2. We have found that the shock speed varied in a broad range. At the start, the shock velocity was relatively higher compared to later days, while it became nearly constant with some fluctuations in later time. Such fluctuations can be due to the limitation of assuming constant velocity of the shock propagation. For a comparison, we refer to Chakrabarti et al. (2005). The authors showed that considering the constant velocity approximation, the average speed of the shock is close to 20 m s-1, which is nearly consistent with our result. This also depends on the non-constant time difference between the exposures. However, the overall velocity profile agrees that the shock moved faster when $X_{s}$ was bigger, decreased with $X_{s}$, and afterward almost with constant velocity.